a 420 gram football is thrown horizontally at 50mph. if the QB takes .2 s to throw the ball, what is the magnitude of the ball just after it is thrown and what is the magnitude of Annette force on the ball during the throw?
Answer #1

Given that,

m = 420 gm = 0.42 kg

v = 50 mph = 22.35 m/s

(a)

Magniude of momentum of bal,

P = mv

P = 0.42*22.35

P = 9.38 kg.m/s

(b)

Change in momentum, deltaP = 9.38 - 0 = 9.38

force on the ball during the throw,

F =  deltaP / t

F = 9.38 / 0.2

F = 46.93 N